OpenStudy (anonymous):

divide and simplify Please help will do the problem on the equation so it looks better

OpenStudy (anonymous):

\[\frac{ 2x ^{2}-6x }{ 5x-1 }\div \frac{ x-3 }{ 5x-1 }\]

OpenStudy (anonymous):

Ok, that equals:\[\frac{ 2x(x - 3) }{ 5x - 1 } \times \frac{ 5x - 1 }{ x - 3 }\]The 5x - 1 factor and the x - 3 factor will cancel, so your answer will be left as: 2x And that's it!

OpenStudy (anonymous):

thanks tcarroll you are awesome, how do you make it look so simple?

OpenStudy (anonymous):

can you help me with some more please

OpenStudy (anonymous):

I'm just lucky. Students like you are awesome to work with!

OpenStudy (anonymous):

Sure, if I can help, I will!

OpenStudy (anonymous):

how about this one \[\frac{ 16a ^{2} }{4a+9b}-\frac{ 81b ^{2} }{ 4a+9b}\]

OpenStudy (anonymous):

I like how people take you step by step for me to understand instead of writing the answer and not letting me know how they got it.

OpenStudy (anonymous):

would you divide the 16 by 4 and the 81 by 9

OpenStudy (anonymous):

\[\frac{ 16a ^{2} - 81b ^{2} }{ 4a + 9b } = \frac{ (4a + 9b) \times (4a - 9b) }{ 4a + 9b }\]which equals: 4a - 9b because we have another canceling of a common term in both the numerator and the denominator

OpenStudy (anonymous):

That common factor was 4a + 9b which cancels out, leaving as an answer: 4a - 9b And I'm glad you recognize that I try to give explanations. That means that you are a good, hard-working student who wants to learn. That is very commendable because too many students here just want answers and sadly, they will not learn that way.

OpenStudy (anonymous):

i had 4a+9b but now I see how you got the 4a-9b. k please hold on

OpenStudy (anonymous):

ok next question is LCM(least common multiple) x^2+7x,x^2-7x

OpenStudy (anonymous):

would it b 7

OpenStudy (anonymous):

no holf on it would be 0

OpenStudy (anonymous):

Ok, the key to LCM problems is first to factor what you have. So, what you really have to start out with is: x(x + 7) and x(x - 7) So, you have to include as factors: x, x + 7, and x - 7 So, your answer is: x(x + 7)(x - 7) which is x^3 - 49x or x(x^2 - 49) You only need one "x" because each term has an "x" in it.

OpenStudy (anonymous):

After I asked you if it was 7 or 0 I remeber that I had a problem like this written down from the past I cam up with x(x^2-49)

OpenStudy (anonymous):

Yes, that's good that you independently came up with the right answer! That matches mine and you got it right. Good job!

OpenStudy (anonymous):

ok here is another one, I wish that I could get the concepts of these a lot easier like you do here is a new one that i have to simplify expression to lowest term and then leave in factor form \[\frac{ 3y }{ y(4y-1) }+\frac{ 2 }{ 4y-1}\]

OpenStudy (anonymous):

Ok, and nice working with you, but I only have time for this last one.\[\frac{ 3 }{ 4y - 1 } + \frac{ 2 }{ 4y - 1 } = \frac{ 5 }{ 4y - 1 }\]Notice how the "y" canceled out of the first term before we added them.

OpenStudy (anonymous):

yes and I got 5/4y-1 is that the correct answer

OpenStudy (anonymous):

Yes, you are getting these well! Good work and it was very nice working with you!

OpenStudy (anonymous):

you too thanks so much I am doing my final and wanting to make sure that I am doing it correctly. This is my last class for my aa, I have a a in this class but I just want to make sure that I am doing everything right before I answer thanks so much again

OpenStudy (anonymous):

You'll do well because you are a good student and a nice person. Both count.